OpenStudy (anonymous):

How would you graph 4x + 3y – 9 = 0?

OpenStudy (kanwal32):

ok

OpenStudy (anonymous):

4x + 3y - 9 = 0 4x - 4x + 3y - 9 - 9 = -4x + 9 3y = -4x + 9

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

3y/3 = -4x/3 + 9/3 y = (-4/3)x -3 x = all real numbers y = all real numbers

OpenStudy (anonymous):

welcome :)